OpenStudy (anonymous):

Find (f+g)(x),(f-g)(x), (fg)(x), and (f/g)(x), and their domains. f(x)=x+6, g(x)=x-6 (f+g)(x)=??

sam (.sam.):

f(x)=x+6, g(x)=x-6 (f+g)(x)=(x+6)+(x-6)

OpenStudy (anonymous):

So the values are literally x+6 and x-6 I keep thinking I need to solve for something??

OpenStudy (anonymous):

So like...(f-g) is just (x+6)-(x-6)??

sam (.sam.):

\[(f+g)(x)=(x+6)+(x-6)\] \[(f-g)(x)=(x+6)-(x-6)\] \[(fg)(x)=(x+6)(x-6)\] \[(f/g)(x)=\frac{(x+6)}{(x-6)}\] Can you see the pattern here?

OpenStudy (anonymous):

Yeah...I always think I am solving for x

sam (.sam.):

No, you're not solving here

OpenStudy (anonymous):

I got it, rather simple andd I'm overthinking it. It helps to have someone confirm it. Thanks again!